OpenStudy (henrietepurina):

Algebra 1 question :D Medal + Fan!

OpenStudy (henrietepurina):

Matt sells burgers and sandwiches. The daily cost of making burgers is $520 more than the difference between the square of the number of burgers sold and 30 times the number of burgers sold. The daily cost of making sandwiches is modeled by the following equation: C(x) = 2x^2 - 40x + 300 C(x) is the cost in dollars of selling x sandwiches. Which statement best compares the minimum daily cost of making burgers and sandwiches? It is greater for sandwiches than burgers because the approximate minimum cost is $250 for burgers and $292 for sandwiches. It is greater for sandwiches than burgers because the approximate minimum cost is $100 for burgers and $295 for sandwiches. It is greater for burgers than sandwiches because the approximate minimum cost is $295 for burgers and $100 for sandwiches. It is greater for burgers than sandwiches because the approximate minimum cost is $292 for burgers and $250 for sandwiches.

OpenStudy (henrietepurina):

@jagr2713 mind helping me?

OpenStudy (henrietepurina):

hey @surjithayer :D

jagr2713 (jagr2713):

The cost of making 0 sandwiches is C(0)=2(0)2−40(0)+300= The cost of making 0 burgers is B(0)=(0)2−30(0)+520=

OpenStudy (henrietepurina):

hi y'all :D

jagr2713 (jagr2713):

so it's 300 and 520 but that is not one of my options

OpenStudy (henrietepurina):

yeah.. weird?!

jagr2713 (jagr2713):

how D;

OpenStudy (misty1212):

\[x^2-30x+520\] is how i read it

jagr2713 (jagr2713):

me to

OpenStudy (misty1212):

and \[C(x) = 2x^2 - 40x + 300\]

OpenStudy (henrietepurina):

no way? let me try and see if I made a mistake in the writing ok?

jagr2713 (jagr2713):

@misty1212 when did u start OS

OpenStudy (misty1212):

and your job is to find the minimum, which is the same as the vertex

OpenStudy (henrietepurina):

what if the minimum cost is for 1 of each?

OpenStudy (henrietepurina):

not 0?

OpenStudy (misty1212):

hold the phone dear, we are not nearly there yet we have more work to do

OpenStudy (henrietepurina):

ok ok :|

OpenStudy (misty1212):

@jagr2713 idk maybe three weeks?

OpenStudy (misty1212):

to find the minimum value of a quadratic the first coordinate is \(-\frac{b}{2a}\)

jagr2713 (jagr2713):

i was wondering b.c ur everywhere lol i would see a question and start typing and ur done already and i am like wow lol :D u work fast

OpenStudy (henrietepurina):

haha... she's good alright :D

OpenStudy (misty1212):

for \[C(x) = 2x^2 - 40x + 300\] \(a=2,b=-40\) so \(-\frac{b}{2a}=\frac{40}{4}=10\)

OpenStudy (misty1212):

find \(C(10)\) use a calculator

OpenStudy (misty1212):

for \[x^2-30x+520\] \(a=1,b=-30,-\frac{b}{2a}=\frac{30}{2}=15\)

OpenStudy (misty1212):

so you can compare them

OpenStudy (henrietepurina):

if C(10) then the equation equals -60

OpenStudy (misty1212):

hmmm no

OpenStudy (misty1212):

try again

OpenStudy (henrietepurina):

ok

OpenStudy (henrietepurina):

I keep getting -60, mind helping?

OpenStudy (henrietepurina):

how? 100.... what I am I doing wrong?

OpenStudy (misty1212):

messing up inputting it in to the calculator is my guess your calculator cannot read your mind, use parentheses if needed

OpenStudy (misty1212):

or just do the arithmetic without one

OpenStudy (henrietepurina):

yeah... that must be it :)

OpenStudy (henrietepurina):

so would the answer be C then?

OpenStudy (misty1212):

and now we are done, because there is only one choice where the cost of sandwiches is $100

OpenStudy (misty1212):

it is always C so it is probably C now too

OpenStudy (henrietepurina):

haha... yeah :D thank you so much!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (henrietepurina):

nice heart :D

OpenStudy (henrietepurina):

what could @surjithayer be typing for so long? :P

OpenStudy (anonymous):

\[B(x)=x^2-3ox+520\] \[B'(x)=2x-30\] B'(x)=0 gives 2 x-30=0,x=15 B"(x)=2>0 at x=15 cost of burger is minimum at x=15 or minimum cost of burgers=(15)^2-30*15+520=225-450+520=295 c(x)=2x^2-40x+300 c'(x)=4x-40 c'(x)=0 gives 4x-40=0,x=10 c"(x)=4>0 at x=10 cost of sandwiches is minimum at x=10 and minimum cost of sandwiches=2(10)^2-40*10+300=200-400+300=100